# Hallucination Detection in Foundation Models for Decision-Making: A Flexible Definition and Review of the State of the Art

### v1

@@ 2.2 Robotics
Foundation models have also been used in the robotics domain for object detection, affordance prediction, grounding, navigation, and communication. Ichter et al. (2023) are motivated by the issue of misalignment between the capabilities of a robot and what an LLM believes it is capable of performing. Because LLMs may not specifically be trained with data from the robot it is to be deployed on, there is a gap in the model’s understanding and the true capacity of the robot, which could lead to hallucinated generations that cannot feasibly be used at runtime. The authors propose SayCan as a method to combine the general knowledge of LLMs with the specific capabilities of a robot in the real-world. Specifically, an LLM is given a task in text form, and is asked to output a list of smaller actions to take in order to complete said task successfully. To constrain the LLM to generate possible actions available to the robot, they assume access to (1) the probability distribution of next tokens to generate from the model, and (2) a set of available skills on the robot, with which they compute the probability of the LLM generating each of the skills next. SayCan greedily selects the action that has the highest product of the next token probability from the LLM and the probability of the action actually successfully being executed in the environment, until the model predicts it has completed the task.
@@
Rather than relying purely on textual context, PaLM-E, proposed by Driess et al. (2023), is a multi-modal model that converts various sensor inputs (e.g., images) to a token-space embedding that is combined with instruction embeddings to be input to a PaLM LLM Chowdhery et al. (2023). PaLM is used to either answer questions about the surroundings of the robot, or to plan a sequence of actions to perform to complete a task. Driess et al. (2023) further acknowledge that the multi-modality of their PaLM-E architecture leads to increased risk of hallucinations.
@@
Inspired by recent promising findings in using foundation models to generate programs Chen et al. (2021), other works deploy foundation models to write low-level code to be run on robots. Liang et al. (2023) present Code as Policies, which uses LLMs to hierarchically generate interactive code and functions that can be called. As the model writes main code to be run on a robot given an instructive prompt of the task from the user, it identifies functions to call within the higher level code to complete the task successfully. The authors show that LLMs can leverage third party libraries for existing functions, or develop their own library of functions dynamically with custom methods for the task. While the functionality of Code as Policies can be tested easily for low-level skill definitions, longer multi-step problems require testing whether all requested conditions have been met by running the generated code on the robot. As such, Hu et al. (2024) propose the RoboEval performance benchmark for testing robot-agnostic LLM-generated code. Specifically, the CodeBotler platform provides an LLM access to abstract functions like “pick,” “place,” and “get_current_location” that have the same external interface regardless of the robot to be deployed on. Like Code as Policies, CodeBotler is provided a text instruction from the user and generates code to be tested. Then the RoboEval benchmark uses RoboEval Temporal Logic (RTL) to test whether the generated code meets task and temporal ordering constraints provided in the original prompt. Furthermore, they test the robustness of the LLM by passing in several paraphrased prompts to check for consistency across inputs. We discuss similar consistency-checking strategies for identifying hallucinations in decision-making tasks further in Section 4.3.1.
@@
In the space of robot navigation, LM-Nav leverages a VLM and attempts to predict a sequence of waypoints for a robot to follow and visit landmarks described within a language command Shah et al. (2023). Here, the authors use in-context learning Dong et al. (2023) to teach GPT-3 to extract desired landmarks from a natural language instruction. Assuming there are images of the possible landmarks the robot can navigate to in its environment, LM-Nav uses CLIP Radford et al. (2021) to predict the closest matching pairs of extracted landmark descriptions and waypoint images. Finally, dynamic programming is applied on the complete graph of the environment to optimize the path of landmarks to visit. The overall predicted path is optimized to maximize the likelihood of successfully completing the instruction input to the model.
2.3 Other Areas
There are also other works that apply foundation models for decision-making outside of the robotics and autonomous vehicle domains. For example, ReAct from Yao et al. (2023b) identifies that a key limitation of chain-of-thought reasoning Wei et al. (2022) is that the model does not update its context or action based on observations from an environment. As such, chain-of-thought reasoning relies purely on the internal reasoning of the foundation model itself to predict actions to take, missing a crucial step in grounding its actions with their effects on the environment. Given a prompt, ReAct iterates between an internal reasoning step and acting in the environment to build up context relevant to the task. Yao et al. (2023b) showcase the promise of the method in a QA setting where the LLM can take actions to query information from an external knowledge base, as well as an interactive text-based game, ALFWorld Shridhar et al. (2021). Chen et al. (2023b) admit that ReAct is a powerful tool for dynamic reasoning and grounding, but is limited by the fact that the updated context from the Act step is only helpful for the particular task the model is currently deployed for. They propose Introspective Tips to allow an LLM to reason about its past successes and failures in a world to generate general tips that will be helpful across diverse instruction-following tasks. Specifically, tips are generated from the past experience of the model from a similar set of tasks, from expert demonstrations, and from several games that differ from the target task. By summarizing these experiences into more concise tips, Chen et al. (2023b) show that Introspective Tips outperform other methods in ALFWorld with both few- and zero-shot contexts.
Park et al. (2023) and Wang et al. (2023b) apply foundation models in more complex environments to push models to their limits to simulate realistic human behaviors and test lifelong learning. Park et al. (2023) propose generative agents that produce believable, human-like interactions and decisions within a small town sandbox environment. They develop a module for individual agents in the simulation to store and retrieve memories, reflect about past and current experiences, and interact with other agents. Their generative agents use similar methods to ReAct and Introspective Tips to act based on a memory of experiences, but also interact and build relationships with other agents through dialogue. The authors show that the agents are able to effectively spread information, recall what has been said to others and stay consistent in future dialogue interactions, and coordinate events together. Sometimes, however, agents are found to hallucinate and embellish their responses with irrelevant details that may be attributed to the training dataset of outside, real-world knowledge. Voyager, from Wang et al. (2023b), deploys GPT-4 to the MineDojo environment Fan et al. (2022) to test its in-context lifelong learning capabilities. The architecture prompts GPT-4 to generate next high-level tasks to complete, given the agent’s current state and results of past tasks — a form of automatic curriculum generation. Voyager then identifies what intermediate general skills would be required to complete the task, and the LLM is used to fill in a skill library with helpful low-level skills in the form of programs that call functions that are available to the simulator. GPT-4 is prompted to generate skills that are generalizable to multiple tasks, so that the skill generation step does not have to be called for every task if the skill is already stored in the library. Wang et al. (2023b) show that Voyager continuously learns to explore the diverse tech tree available within MineDojo while building and leveraging skills. Even so, they find that the LLM hallucinates when generating tasks to tackle and when writing the code to execute for a particular skill, discussed further in Section 3.2.
Kwon et al. (2023) explore the use of LLMs to act as a proxy for a hand-tuned reward function in RL tasks. This application is particularly motivated by decision-making tasks that are difficult to specify with a reward function, but can be explained textually with preferences of how a policy should generally act. Specifically, the LLM evaluator first undergoes in-context learning with examples of how it should decide the reward in several cases of the task that the agent will be deployed to. Then, during RL training, the LLM is provided a prompt with the trajectory of the agent within the episode, the resulting state from the simulator, and the original task objective from the user, and is asked to generate a binary reward for the agent (1 if success, 0 else). The binary reward is added to the experience replay, and the agent can be updated using any RL algorithm. Kwon et al. (2023) find that a baseline in their work that predicts rewards with no in-context learning especially hallucinates with incoherent reasoning.
3 Hallucinations
@@ 3.1 What are hallucinations?
Across current literature on foundation models, there exist similar patterns and themes that can be used to develop a unified definition for hallucinations. With the majority of works studying this problem within QA tasks, where ground truth answers are available, several authors explain hallucinations as producing text that includes details/facts/claims that are fictional/misleading/fabricated rather than truthful or reliable Rawte et al. (2023). Works making use of a dedicated knowledge-base further describe hallucinations as generating nonsensical or false claims that are unsubstantiated or incorrectly cited Mündler et al. (2024); Chen et al. (2023a); Zhang et al. (2023b); Li et al. (2023b). Varshney et al. (2023) also present the idea that foundation models may sound syntactically correct, or coherent, while simultaneously being incorrect. Gallifant et al. (2024), who perform a peer review of the GPT-4 technical paper, state that hallucinations include responses that are irrelevant to the original prompt. Li et al. (2023c), who specifically explore hallucinations of LVLMs in detecting and classifying objects within images, define hallucinations as generating object descriptions inconsistent with target images. A common theme among existing hallucination definitions for QA, information retrieval, and image captioning domains is that, while the generation may sound coherent, either the output is incorrect, or the model’s reasoning behind the generated text is incorrect. However, we find these characteristics on their own do not completely encompass the hallucinations found in decision-making tasks in literature, thus requiring additional nuances.
@@
Within papers that apply foundation models to decision-making tasks specifically, researchers have encountered similar problems of hallucinations impacting performance. Park et al. (2024) describe hallucinations as predicting an incorrect feasibility of an autonomous system when generating an explanation behind the uncertainty of an action to take. Similarly, Kwon et al. (2023) find that language models may provide incoherent reasoning behind their actions. Wang et al. (2024) and Ren et al. (2023) believe that these generative models also have a sense of high (false) confidence when generating incorrect or unreasonable plans. In the case of robot navigation and object manipulation, Hu et al. (2024) and Liang et al. (2024) refer to hallucinations as attempting to interact with non-existent locations or objects.
Table 1: Examples of applying Definition 3.1 to different tasks. Note that developers may choose to only define a subset of hallucination metrics for their deployment depending on evaluation preferences.
Metric
Problem Setting
Questing-Answering
Image Captioning
Planning
Control
Consistency
Generations must align with database facts
Objects in description must appear in image
Predicted sub-task must be feasible to solve
Predicted action must be possible to perform
Desired Behavior
Tone of answer should be informative
Censor descriptions for inappropriate images
Plans should maximize expected return
Predict actions to complete plan efficiently
Relevancy
Answers should not include references to unrelated topics
Descriptions should not be embellished with details that cannot be confirmed
Predicted sub-tasks and actions should not stray from the end goal with unnecessary steps
Plausibility
Generation is syntactically sound and believable
Generated plan is reasonable and seems to attempt to accomplish goal
@@
In the code generation task, Chen et al. (2021) use the term “alignment failure,” with similar effects to those of hallucinations discussed above. More specifically, the authors informally describe an alignment failure as an outcome where a model is capable of performing a task, but chooses not to. If a model is able to complete a task successfully within its latent space (perhaps through additional prompt engineering or fine-tuning), one may ask, “Why would the model choose not to?” As foundation models are trained with the next-token reconstruction objective on a training set, they attempt to maximize the likelihood of the next token appearing at test-time as well. Consequently, if the test-time prompt includes even minor mistakes, Chen et al. (2021) find that LLMs will continue to generate buggy code to match the input prompt. This issue is further described in Section 3.3.
@@
We realize existing definitions for hallucinations are extremely disparate depending on the deployment area. As such, to bridge existing QA application areas, decision-making tasks, and all other possible test scenarios for foundation models, we combine these findings and define the term hallucination as follows:
@@
    • A list of constraints that must be consistent within the generation.
@@
  3. 3.
    Metrics measuring consistency, desirability, relevancy, and syntactic soundness (plausibility) of generations.
In practice, this definition generally encapsulates the qualities of hallucinations discussed earlier. For example, in QA or object detection tasks, one may define a set of relevant topics that a generation should not stray from, and constraints may be held in the form a knowledge-base of ground truth facts. The desired behavior of the generation may be to be phrased in an informative manner, rather than sarcastic. On the other hand, in robot manipulation settings, a developer may have a set of constrained actions feasible on the robot, and the desired behavior could be to complete a task with as few actions as possible. Relevancy may be measured in relation to the specific task to be deployed on (e.g., a prompt requesting a recipe to make pasta would find it irrelevant if the model also suggested a song to play while cooking). Finally, plausibility informally relates to a measure of how believable an output is to a critic. A more realistic generation has a greater chance of deceiving the user into trusting the model, even when the plan may be hallucinated. Overall, hallucinated outputs may contain one or more of the core characteristics (inconsistent, undesired, irrelevant, and plausible) simultaneously, and our definition can be flexibly applied to any deployment scenario in mind by choosing metrics for each characteristic, respectively. We show more examples of applying our definition to various tasks in Table 1.
3.2 Examples
@@
As discussed in Section 2.1Wen et al. (2023) test GPT-4V on the autonomous driving task and identify failure modes. Regardless of the weather and driving conditions, GPT-4V has difficulty detecting and identifying the traffic light state at an intersection, until the image has zoomed in on the light itself. It also presents additional irrelevant (or completely false) details about other agents, when the prompt had no mention of them in the first place. Furthermore, the model also has difficulty in describing temporal sequences (i.e., videos) and categorizing images by their direction within a panoramic view from the vehicle’s perspective. In their later work, Wen et al. (2024) describe that hallucinations arise in these complex environments because of the high variability in driving scenarios. Even after applying hallucination mitigation techniques like chain-of-thought reasoning, the model is not free of these undesired outputs. A similar work evaluating the frequency at which LVLMs hallucinate in their descriptions of images, finds that these models’ outputs may include non-existent objects, or additional irrelevant phrases (that may not even be possible to test for accuracy) Li et al. (2023c). For example, in a picture of food on a table, an LVLM hallucinates a non-existent beverage, and predicts that the “table is neatly arranged, showcasing the different food items in an appetizing manner.” Although the classification error and irrelevant generation in this example are not critical, earlier works warn of possible failures with more severe, high societal impact (e.g., biases in models leading to marginalizing users) Bommasani et al. (2022).
@@
Chen et al. (2021) explore alignment failures of LLMs applied to code completion tasks. The authors evaluate the likelihood of these models generating defective code given different input prompts, and discover that in-context learning using examples with buggy code has a higher chance of resulting in poor generations from the model on the actual task at hand. The study also identifies similar model biases towards race, gender, religion, and other representations. Furthermore, the authors find that their model, Codex, is able to generate code that could assist with developing insecure applications or malware, albeit in a limited manner. These findings have been corroborated by other foundation model code generation works in the robotics domain. For example, Wang et al. (2023b) describe that Voyager sometimes generates code with references to items that do not exist within MineDojo. Similarly, Hu et al. (2024) find that their model has the tendency to call functions with invalid objects or locations, pickup objects when it is already holding something, ask for help when no one is near, and other undesired behaviors.
@@
Several works focus on identifying cases of hallucinations in QA tasks. Although this application area is not the direct focus of this work, we present examples of hallucinations in this field as we can glean similar failure modes that could arise within decision-making systems. Common hallucinations in QA result in incorrect answers to questions. For example, Achiam et al. (2023) find that GPT-4 “hallucinates facts and makes reasoning errors.” Achiam et al. (2023) categorize these failures into closed-domain (given context, the model generates irrelevant information that was not in the context) and open-domain (the model outputs incorrect claims without any context) hallucinations. After fine-tuning on more data with a hallucination mitigation objective, the model reduces its tendency to hallucinate, but still does not achieve perfect accuracy — a similar trend encountered by Touvron et al. (2023a). Another set of works identify hallucinations with contradictions among several sampled generations from an LLM, discussed further in Section 4.3.1 Mündler et al. (2024); Zhang et al. (2023b). Intuitively, if a context passed into a model results in conflicting generations, the model must be hallucinating some part of the output. Notice in this example, with relation to Definition 3.1, self-contradiction works test for consistency among multiple (hallucinated) generations, rather than with respect to a ground-truth knowledge-base that usually exists in QA tasks. As such, our definition can flexibly apply to different system setups by describing consistency, desired behavior, and relevancy respectively.
3.3 Why do they happen?

### v2

@@ 2.2. Robotics
Foundation models have also been used in the robotics domain for object detection, affordance prediction, grounding, navigation, and communication. An example of a robot deployed with LVLM capabilities and potential hallucinations is shown in Figure 3 of Appendix A.1. Ichter et al. (2023) are motivated by the issue of misalignment between the capabilities of a robot and what an LLM believes it is capable of performing. Because LLMs may not specifically be trained with data from the robot it is to be deployed on, there is a gap in the model’s understanding and the true capacity of the robot, which could lead to hallucinated generations that cannot feasibly be used at runtime. The authors propose SayCan as a method to combine the general knowledge of LLMs with the specific capabilities of a robot in the real-world. Specifically, an LLM is given a task in text form, and is asked to output a list of smaller actions to take in order to complete said task successfully. To constrain the LLM to generate possible actions available to the robot, they assume access to (1) the probability distribution of next tokens to generate from the model, and (2) a set of available skills on the robot, with which they compute the probability of the LLM generating each of the skills next. SayCan greedily selects the action that has the highest product of the next token probability from the LLM and the probability of the action actually successfully being executed in the environment, until the model predicts it has completed the task.
@@
Rather than relying purely on textual context, PaLM-E, proposed by Driess et al. (2023), is a multi-modal model that converts various sensor inputs (e.g., images) to a token-space embedding that is combined with instruction embeddings to be input to a PaLM LLM (Chowdhery et al., 2023). PaLM is used to either answer questions about the surroundings of the robot, or to plan a sequence of actions to perform to complete a task. Driess et al. further acknowledge that the multi-modality of their PaLM-E architecture leads to increased risk of hallucinations.
@@
Inspired by recent promising findings in using foundation models to generate programs (Chen et al., 2021), other works deploy foundation models to write low-level code to be run on robots. Liang et al. (2023) present Code as Policies, which uses LLMs to hierarchically generate interactive code and functions that can be called. As the model writes main code to be run on a robot given an instructive prompt of the task from the user, it identifies functions to call within the higher level code to complete the task successfully. The authors show that LLMs can leverage third party libraries for existing functions, or develop their own library of functions dynamically with custom methods for the task. While the functionality of Code as Policies can be tested easily for low-level skill definitions, longer multi-step problems require testing whether all requested conditions have been met by running the generated code on the robot. As such, Hu et al. (2024) propose the RoboEval performance benchmark for testing robot-agnostic LLM-generated code. Specifically, the CodeBotler platform provides an LLM access to abstract functions like “pick,” “place,” and “get_current_location” that have the same external interface regardless of the robot to be deployed on. Like Code as Policies, CodeBotler is provided a text instruction from the user and generates code to be tested. Then the RoboEval benchmark uses RoboEval Temporal Logic (RTL) to test whether the generated code meets task and temporal ordering constraints provided in the original prompt. Furthermore, they test the robustness of the LLM by passing in several paraphrased prompts to check for consistency across inputs. We discuss similar consistency-checking strategies for identifying hallucinations in decision-making tasks further in Section 4.3.1.
@@
In the space of robot navigation, LM-Nav leverages a VLM and attempts to predict a sequence of waypoints for a robot to follow and visit landmarks described within a language command (Shah et al., 2023). Here, the authors use in-context learning (Dong et al., 2023) to teach GPT-3 (Brown et al., 2020) to extract desired landmarks from a natural language instruction. Assuming there are images of the possible landmarks the robot can navigate to in its environment, LM-Nav uses CLIP (Radford et al., 2021) to predict the closest matching pairs of extracted landmark descriptions and waypoint images. Finally, dynamic programming is applied on the complete graph of the environment to optimize the path of landmarks to visit. The overall predicted path is optimized to maximize the likelihood of successfully completing the instruction input to the model.
3. Hallucinations
@@ 3.1. What are hallucinations?
Across current literature on foundation models, there exist similar patterns and themes that can be used to develop a unified definition for hallucinations. With the majority of works studying this problem within QA tasks, where ground truth answers are available, several authors explain hallucinations as producing text that includes details/facts/claims that are fictional/misleading/fabricated rather than truthful or reliable (Rawte et al., 2023). Works making use of a dedicated knowledge-base further describe hallucinations as generating nonsensical or false claims that are unsubstantiated or incorrectly cited (Mündler et al., 2024; Chen et al., 2023b; Zhang et al., 2024b; Li et al., 2023a). Varshney et al. (2023) also present the idea that foundation models may sound syntactically correct, or coherent, while simultaneously being incorrect. Gallifant et al. (2024), who perform a peer review of the GPT-4 technical paper, state that hallucinations include responses that are irrelevant to the original prompt. Li et al. (2023b), who specifically explore hallucinations of LVLMs in detecting and classifying objects within images, define hallucinations as generating object descriptions inconsistent with target images. A common theme among existing hallucination definitions for QA, information retrieval, and image captioning domains is that, while the generation may sound coherent, either the output is incorrect, or the model’s reasoning behind the generated text is incorrect. However, we find these characteristics on their own do not completely encompass the hallucinations found in decision-making tasks in literature, thus requiring additional nuances.
@@
Within papers that apply foundation models to decision-making tasks specifically, researchers have encountered similar problems of hallucinations impacting performance. Park et al. (2024) describe hallucinations as predicting an incorrect feasibility of an autonomous system when generating an explanation behind the uncertainty of an action to take. Similarly, Kwon et al. (2023) find that language models may provide incoherent reasoning behind their actions. Wang et al. (2024b) and Ren et al. (2023) believe that these generative models also have a sense of high (false) confidence when generating incorrect or unreasonable plans. In the case of robot navigation and object manipulation, Hu et al. (2024) and Liang et al. (2024) refer to hallucinations as attempting to interact with non-existent locations or objects.
@@
In the code generation task, Chen et al. (2021) use the term “alignment failure,” with similar effects to those of hallucinations discussed above. More specifically, the authors informally describe an alignment failure as an outcome where a model is capable of performing a task, but chooses not to. If a model is able to complete a task successfully within its latent space (perhaps through additional prompt engineering or fine-tuning), one may ask, “Why would the model choose not to?” As foundation models are trained with the next-token reconstruction objective on a training set, they attempt to maximize the likelihood of the next token appearing at test-time as well. Consequently, if the test-time prompt includes even minor mistakes, Chen et al. find that LLMs will continue to generate buggy code to match the input prompt. This issue is further described in Section 3.3.
We realize existing definitions for hallucinations are extremely disparate depending on the deployment area. Nevertheless, we have identified three distinct features that are commonly missing within hallucinated generations: compliancedesirability, and relevancy. Additionally, we find hallucinations may seem to be plausible while they are in fact unacceptable. A compliance metric checks that the generation meets hard constraints while the desirability metric measures how well the generation meets soft constraints defined by the model engineer. Irrelevant generations refer to predictions that contain details outside of a set of requested topics. Notice that irrelevant predictions can still be considered compliant and desirable depending on how the hard and soft constraints are defined. Plausibility compares the generation’s syntax to those of a set of known, unhallucinated samples. While relevancy is evaluating the content of a prediction, plausibility is assessing its phrasing. Table 1 provides complete definitions for each of the four characteristics.
Table 1. Definitions for compliance, desirability, relevancy, and plausibility — the four characteristics of hallucinations.
Characteristic Definition
Compliance Generation meets hard constraints that cannot be ignored
Desirability Generation attempts to meet soft constraints measured by some cost or reward function
Relevancy Contents of generation do not fall outside of a defined set of topics
Plausibility The syntactic similarity of a generation and in-domain normal samples measured via a critic function
@@
Then, to bridge definitions from existing QA application areas, decision-making tasks, and all other possible test scenarios for foundation models, we combine these findings and define the term hallucination as follows:
@@
    • A list of constraints that must be compliant within the generation.
@@
  3. (3)
    Metrics measuring compliance, desirability, relevancy, and syntactic soundness (plausibility) of generations.
In practice, this definition generally encapsulates the qualities of hallucinations discussed earlier. For example, in QA or image captioning tasks, one may define a set of relevant topics that a generation should not stray from, and constraints may be held in the form a knowledge-base of ground truth facts. The desired behavior of the generation may be to be phrased in an informative manner, rather than sarcastic. On the other hand, in robot manipulation settings, a developer may have a set of constrained actions feasible on the robot, and the desired behavior could be to complete a task with as few actions as possible. Relevancy may be measured in relation to the specific task to be deployed on (e.g., a prompt requesting a recipe to make pasta would find it irrelevant if the model also suggested a song to play while cooking). Finally, plausibility informally relates to a measure of how believable an output is to a critic. A more realistic generation has a greater chance of deceiving the user into trusting the model, even when the plan may be hallucinated. Overall, hallucinated outputs may contain one or more of the core characteristics (noncompliant, undesired, irrelevant, and plausible) simultaneously, and our definition can be flexibly applied to any deployment scenario in mind by choosing metrics for each characteristic, respectively. We show more examples of applying our definition to various tasks in Table 2.
3.2. Examples
Table 2. Examples of applying Definition 3.1 to different tasks. Note that developers may choose to only define a subset of hallucination characteristics for their deployment depending on evaluation preferences. The table is split into non-decision-making and decision-making applications.
Problem Setting Characteristic
Compliance Desired Behavior Relevancy Plausibility
Question-Answering Generations must align with database facts Tone of answer should be informative Answers should not include references to unrelated topics Generation is syntactically sound and believable
Image Captioning Objects in description must appear in image Censor descriptions for inappropriate images Descriptions should not be embellished with details that cannot be confirmed
Planning Predicted sub-task must be feasible to solve Plans should maximize expected return Predicted sub-tasks and actions should not stray from the end goal with added steps Generated plan is reasonable and seems to attempt to accomplish goal
Control Predicted action must be possible to perform Predict actions to complete plan efficiently
@@ 3.2. Examples
As discussed in Section 2.1Wen et al. (2024b) test GPT-4V on the autonomous driving task and identify failure modes. Regardless of the weather and driving conditions, GPT-4V has difficulty detecting and identifying the traffic light state at an intersection, until the image has zoomed in on the light itself. It also presents additional irrelevant (or completely false) details about other agents, when the prompt had no mention of them in the first place. Furthermore, the model also has difficulty in describing temporal sequences (i.e., videos) and categorizing images by their direction within a panoramic view from the vehicle’s perspective. In their later work, Wen et al. (2024a) describe that hallucinations arise in these complex environments because of the high variability in driving scenarios. Even after applying hallucination mitigation techniques like chain-of-thought reasoning, the model is not free of these undesired outputs. A similar work evaluating the frequency at which LVLMs hallucinate in their descriptions of images, finds that these models’ outputs may include non-existent objects, or additional irrelevant phrases (that may not even be possible to test for accuracy) (Li et al., 2023b). For example, in a picture of food on a table, an LVLM hallucinates a non-existent beverage, and predicts that the “table is neatly arranged, showcasing the different food items in an appetizing manner.” Although the classification error and irrelevant generation in this example are not critical, earlier works warn of possible failures with more severe, high societal impact (e.g., biases in models leading to marginalizing users) (Bommasani et al., 2022).
@@
Chen et al. (2021) explore alignment failures of LLMs applied to code completion tasks. The authors evaluate the likelihood of these models generating defective code given different input prompts, and discover that in-context learning using examples with buggy code has a higher chance of resulting in poor generations from the model on the actual task at hand. The study also identifies similar model biases towards race, gender, religion, and other representations. Furthermore, the authors find that their model, Codex, is able to generate code that could assist with developing insecure applications or malware, albeit in a limited manner. These findings have been corroborated by other foundation model code generation works in the robotics domain. For example, Wang et al. (2024c) describe that Voyager sometimes generates code with references to items that do not exist within MineDojo. Similarly, Hu et al. (2024) find that their model has the tendency to call functions with invalid objects or locations, pickup objects when it is already holding something, ask for help when no one is near, and other undesired behaviors.
@@
Several works focus on identifying cases of hallucinations in QA tasks. Although this application area is not the direct focus of this work, we present examples of hallucinations in this field as we can glean similar failure modes that could arise within decision-making systems. Common hallucinations in QA result in incorrect answers to questions. For example, Achiam et al. (2023) find that GPT-4 “hallucinates facts and makes reasoning errors.” Achiam et al. categorize these failures into closed-domain (given context, the model generates irrelevant information that was not in the context) and open-domain (the model outputs incorrect claims without any context) hallucinations. After fine-tuning on more data with a hallucination mitigation objective, the model reduces its tendency to hallucinate, but still does not achieve perfect accuracy — a similar trend encountered by Touvron et al. (2023a). Another set of works identify hallucinations with contradictions among several sampled generations from an LLM, discussed further in Section 4.3.1 (Mündler et al., 2024; Zhang et al., 2024b). Intuitively, if a context passed into a model results in conflicting generations, the model must be hallucinating some part of the output. Notice in this example, with relation to Definition 3.1, self-contradiction works test for compliance by checking consistency among multiple (hallucinated) generations, rather than with respect to a ground-truth knowledge-base that usually exists in QA tasks. As such, our definition can flexibly apply to different system setups by describing compliance, desired behavior, and relevancy respectively.
Additional examples of hallucinations encountered in image, video, and $3$D generation methods, and broader impacts faced by medical, legal, and finance industries are discussed in Appendix B.1 and B.2.
3.3. Why do they happen?